OpenStudy (abbycross167):

can someone please help me with this math question? please? I'll fan and give a medal...

OpenStudy (abbycross167):

what is the domain of x^2-x-6 ---------- x^2-9x+18 Please help me...

OpenStudy (welshfella):

first fator the top and bottom - can you do that ?

OpenStudy (welshfella):

* factor

OpenStudy (abbycross167):

so I just watched a video on how to do this, and I think I answered it, but would you mind checking my steps? please?

OpenStudy (welshfella):

OK

OpenStudy (abbycross167):

Steps— X1 = -(-9)-3 9-3 --------- = --------- = 3 2*1 2 X2 = -(-9)+3 12 --------- = --------- = 6 2*1 2 The domain is (3,6)

OpenStudy (welshfella):

to be honest I dont know what thats all about!

OpenStudy (abbycross167):

so the domain isn't (3,6)?

OpenStudy (welshfella):

The domain is all the possible values of x For a polynomial its all real x. But for a fraction like this some values of x are not included.

OpenStudy (welshfella):

if we factor top and bottom we get ( x + 2)(x - 3) ------------ (x - 6)(x - 3) Now as x - 3 is common to top and botommwe get (x + 2) ------ (x - 6)

OpenStudy (welshfella):

x can have any value except the one value which makes the denominator zero . ( dividing by 0 is not possible)

OpenStudy (welshfella):

so x - 6 cannot be 0 - so what value of x makes x - 6 zero?

OpenStudy (abbycross167):

x=6?

OpenStudy (welshfella):

right So the domain is all real numbers except 6

OpenStudy (welshfella):

Look at the graph and you'll see there is a break in it at x = 6. https://www.desmos.com/calculator/kq3pywjqsm

OpenStudy (abbycross167):

so thats it? no 3? just 6?

OpenStudy (abbycross167):

thank you for the help!! I really appreciate it!